OpenStudy (anonymous):

If it costs $12 to buy 5 pounds of grapes, how much will you have to pay to buy 9 pounds?

OpenStudy (anonymous):

@SalvadorV

OpenStudy (anonymous):

Set up a proportion for the following situation

OpenStudy (triciaal):

5 lbs cost $12 I lb cost 12/5 9 lbs = 9 times 1lb = 9 x 12/5

OpenStudy (anonymous):

if 5 pounds is $12 then 1 pound is $12/5 (you divide both by 5 so you know how much is the cost of the one pound)

OpenStudy (anonymous):

you just have to multiply by 8, so it is 9*12/5= 108/5